# Low temperature ferroelectric state in strontium titanate microcrystals using in situ multi-reflection Bragg coherent X-ray diffraction imaging

### v1

@@ III.2.1 Crystal B SrCO3 impurities
To determine whether the void regions are twins or impurities, we analyzed the samples using X-ray diffraction (XRD) and Fourier-transform infrared spectroscopy (FTIR). Both methods showed that SrCO3 were present in the samples. Powder XRD (Fig. 10(a)) indicated the presence of SrCO3 in the STO spectrum, particularly noticeable in the 300 nm microcrystals. The FTIR spectrum revealed characteristic carbonate absorption bands, notably around 867 cm-1, 1067 cm-1, and 1440 cm-1, which are due to CO groups [63].
@@ III.2.3 Crystal B radial strain and rotation tensor magnitudes
The radial average magnitude of the tensor components for crystal B was compared at 50 K and 180 K, as shown in Fig. 14. The radial histogram plots differ noticeably from those of crystal A, illustrating the different effects of strain due to dislocations and that of voids. Unlike in crystal A, both the volumetric strain and the tensor components are concentrated at the center of crystal B, seen both in increased magnitude and heterogeneity at small radius. Despite the differences in magnitude and distribution pattern, there is consistency with crystal A regarding the high- and low-temperature states of STO. The low-temperature state, on average, exhibits a higher magnitude and heterogeneity.
@@ IV Discussion
There is a consistent transition in structure below 50 K for crystal A and crystal B, as shown in the heatmaps in Fig. 4, 5, 12, and 13. there appears to be two dominant structural modes for crystal A: one below 50 K and one above 50 K. Neither the heatmaps for the CXDPs nor the tensor components show a sharp change between 50 K and 100 K, but there is a consistent pattern across all heatmaps, suggesting that these two states are present. Below 50K is where low temperature anomalies start, as discussed in the introduction.
@@
When we use the position of the Bragg peak along the position to calculate the lattice parameter (Fig. 15), we observe that the lattice parameter decreases linearly due to thermal contributions () down to 50 K, where it then flattens out. This observation matches the lattice parameter trends on biaxially strained STO thin films reported by He et al. [22], who reported no change in the lattice parameter as a result of the AFD transition, but noted that the lattice parameter remains relatively flat below 50 K. Loetzsch et al. noted a similar lattice parameter trend for the axis near the surface of bulk STO [21]. This sub-50 K transition has also been noted for the thermal expansion coefficient, as reported by Tsunekawa et al. [9]. For crystal A, the coefficient of thermal expansion is between 100 K and 300 K, and between 10 K and 50 K. This could be due to damping of the polar soft modes reported below 50 K by Yamanaka et al. [13].
@@
We would like to connect the structural change below 50 K, with the ferroelectric behavior reported for strained STO from second-harmonic optical measurements [23]. Li et al. reported the emergence of a ferroelectric phase transition in STO thin films depending on temperature and in-plane strain [23]. Based on thermodynamic phase models [23, 64], STO thin films can become ferroelectric at 0 K outside the strain range of to (depending on Landau free energy coefficients), while considering the competing AFD phase. Although the strain fields here are not epitaxially induced, the strain tensor values, especially near dislocation cores and impurities (), exceed this threshold value. Although dislocations can result in other complex phenomena in STO, such as the interaction of dislocation cores with oxygen vacancies [65], dislocations have been shown to stabilize flexoelectric polarization at dislocation cores experimentally and theoretically [66]. Strain gradients are known to break local centrosymmetry, which can lead to the stabilization of a ferroelectric state in STO. If the magnitude of the strain field exceeds the threshold required for a ferroelectric state [23, 64], the strain fields induced by impurities, in the case of crystal B, could also stabilize local polarization.
@@
Another possibility for stabilization is the surface strain caused by the size of the microcrystals compared to the bulk. The atoms on the surface of the crystals experience a higher surface energy relative to the bulk. To relieve the surface energy, the atoms can distort themselves, thereby creating surface strain. Reconstructions of the microcrystals shown in Fig. 11, and figures 16, 17, and 19 in the appendices have regions of lattice expansion (yellow) on the surface, which lead to surface strains Fig. 6 and 14. The surface tension on these microcrystals could be analogous to the epitaxial strain induced on STO thin films [21], suggesting that the ferroelectric state of strained STO thin films [22, 23, 14] could also be relevant for these microcrystals.
@@
Thus, we believe that the low-temperature state observed in crystals A and B likely corresponds to the formation of a ferroelectric phase, caused by the displacement of Ti ions due to interactions with the surrounding lattice strain field. Although the ferroelectric polarization state has been proposed to be along the direction under compressive biaxial strain, and either along a or in-plane direction under tensile strain [23], the polarization state cannot be observed in the BCDI reconstructions. The ferroelectric phase will induce a polarization that changes the lattice parameter slightly, much like the AFD transition, but here we do not observe any split Bragg peaks. Ferroelectric polarization has been probed with BCDI [35], but the image of the ferroelectric state is not clear due to the dominant strain fields induced by the defects.
@@
We do not observe any noticeable changes in the heatmaps (Fig. 4, 5, 12, and 13) corresponding to the cubic-to-tetragonal AFD phase transition at 105 K, nor do we observe a change in lattice parameter or change of slope (thermal expansion coefficient) in Fig. 15. Interestingly, He et al. reported that the AFD transition can occur as much as 50 K higher for biaxially strained STO thin films compared to bulk STO [22]. Their experiment measured a half-order superlattice Bragg peak associated with TiO6 octahedra rotation, and they observed a gradual decrease in intensity across different film thicknesses. Hence, it could be that the AFD phase transition occurs in crystal A and crystal B, but it cannot be detected from our measured crystal Bragg peaks or the lattice parameter change in Fig. 15. Furthermore, the rotation of the TiO6 octahedra around the crystallographic axis and the ratio for the tetragonal phase is below the strain resolution for BCDI. However, the AFD phase transition could also be influenced by the surface area-to-volume ratio and the chemistry of the microcrystals, which are not explored here.
@@ V Conclusion
This study successfully leverages the capability of MBCDI to resolve the residual strain and rotation tensor components for two STO microcrystals as they are cooled down to cryogenic temperatures. Below 50 K, each microcrystal transitioned to its low-temperature state, exhibiting increased strain magnitude and heterogeneity throughout the entire crystal. We attribute this state below 50 K to a ferroelectric state, which is stabilized by the strain gradients induced by the dislocations in crystal A and the SrCO3 impurities in crystal B. We do not observe the AFD phase transition in the MBCDI analysis, even though it should be present in strained samples [22]; however, it might be suppressed by the ferroelectric phase [12, 13]. Another possibility is the size or chemistry dependence of the hydrothermally synthesized STO that alters the phase transition behavior. Our results are the first study of structural modes in STO microcrystals at low temperatures, revealing new information about structural transitions at the nanoscale.
@@ VI Data availability
The processed diffraction patterns, final reconstructions, and data analysis scripts are publicly available at 10.5281/zenodo.13738180.
@@ Acknowledgements.
We would like to thank Steve Shapiro for enlightening us about the neutron scattering work on STO, and Felix Hofmann for thoughtful discussions about dislocation analysis. Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. The authors acknowledge Diamond Light Source for time on Beamline I16 under proposal number MM30687, MM33417, and MM34617. N.Z. and Z.A. acknowledges support from the NSFC (Grant No. 12161141012) S. H., S. C., H. N., M. N., and H. K. acknowledges support from the National Research Foundation of Korea (Grant NRF-2021R1A3B1077076). Work performed at UCL was supported by EPSRC. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.Appendix A Phase retrieval
The reconstruction process for each CXDP was performed independently for each reflection. The cropped CXDP on the detector was padded with zeros to a size of pixels and binned by a factor of two along the detector plane, resulting in a data size of pixels.
Each reconstruction was initialized with a random guess. A guided phasing approach [67] with 60 individuals and four generations was used, employing a geometric average breeding mode. For each generation and population, a block of 20 error reduction (ER) and 180 hybrid input-output (HIO) iterations, with , was repeated three times. This was followed by 20 ER iterations to finalize the object. The shrinkwrap algorithm [68] with a threshold of 0.1 was used to update the real-space support at each iteration. The values for the Gaussian kernel shrinkwrap smoothing for each generation were , respectively. The best reconstruction was determined using a sharpness criterion, as it is the most appropriate metric for crystals containing defects [69]. From the 30 best reconstructions, candidates with a Pearson correlation coefficient greater than 0.965, relative to the candidate most similar to the others, were averaged to produce the final reconstruction. This process was repeated again, with the reconstructions from the first round now as starting guesses.
The overall average 3D spatial resolution was 63.7 nm for crystal A and 61.2 nm for crystal B. This was determined by differentiating the electron density amplitude across the crystal/environment interface for five linearly independent directions and fitting a Gaussian to each of the profiles. The reported spatial resolution was the average full width at half maximum of the Gaussian profiles, across all measured reflections and temperatures for each crystal.
@@ Appendix B Individual reconstructions
The individual reconstructions for crystal A are shown in Fig. 16. Here, the phase vortices, shown within the purple circles, represented the positions of the dislocation cores [55] and were used to determine the dislocation lines discussed further in Appendix C.
Fig. 17 shows the individual reconstructions for crystal B.

### v4

@@ III.2.1 Crystal B SrCO3 impurities
To determine whether the void regions are twins or impurities, we analyzed the samples using x-ray diffraction (XRD) and Fourier-transform infrared spectroscopy (FTIR). Both methods showed that SrCO3 was present in the samples. Powder XRD (Fig. 10(a)) indicated the presence of SrCO3 in the STO spectrum, particularly noticeable in the 300 nm microcrystals. The FTIR spectrum revealed characteristic carbonate absorption bands, notably around 867, 1067, and 1440 cm-1, which are due to CO groups [68].
@@ III.2.2 Crystal B Pearson correlation coefficient
Similar to the results presented in Sec. III.1.2, we employed Pearson’s correlation coefficient heat maps analysis to identify changes in the CXDPs (Fig. 12) and the tensor components (Fig. 13).
@@ III.2.3 Crystal B radial strain and rotation tensor magnitudes
The radial average magnitude of the tensor components for crystal B was compared at 50 and 180 K, as shown in Fig. 14. The radial histogram plots differ noticeably from those of crystal A, illustrating the different effects of strain due to dislocations and that of voids. Unlike in crystal A, both the volumetric strain and the tensor components are concentrated at the center of crystal B, seen both in increased magnitude and heterogeneity at small radius. Despite the differences in magnitude and distribution pattern, there is consistency with crystal A regarding the high- and low-temperature states of STO. The low-temperature state, on average, exhibits a higher magnitude and heterogeneity.
@@ IV Discussion
There is a consistent transition in structure below 50 K for crystal A and crystal B, as shown in the heat maps in Figs. 4, 5, 12, and 13. There appears to be two dominant structural modes for crystal A: one below 50 and one above 50 K. Neither the heat maps for the CXDPs nor the tensor components show a sharp change between 50 and 100 K, but there is a consistent pattern across all heat maps, suggesting that these two states are present. Below 50 K is where low-temperature transitions start, as discussed in the introduction.
@@
When we use the position of the Bragg peak along the position to calculate the lattice parameter (Fig. 15), we observe that the lattice parameter decreases linearly due to thermal contributions () down to 50 K, where it then flattens out. This observation matches the lattice parameter trends on biaxially strained STO thin films reported by He et al. [22], who reported no change in the lattice parameter as a result of the AFD transition, but noted that the lattice parameter remains relatively flat below 50 K. Loetzsch et al. noted a similar lattice parameter trend for the axis near the surface of bulk STO [21]. This sub-50 K transition has also been noted for the thermal expansion coefficient, as reported by Tsunekawa et al. [9]. For crystal A, the coefficient of thermal expansion is between 100 and 300 K, and between 10 and 50 K. This could be due to damping of the polar soft modes reported below 50 K by Yamanaka et al. [13].
@@
A possibility for the structural change below 50 K is ferroelectricity. Ferroelectric behavior has been reported for strained STO from second-harmonic optical measurements [23]. Li et al. reported the emergence of a ferroelectric phase transition in STO thin films depending on temperature and in-plane strain [23]. Based on thermodynamic phase models [23, 69], STO thin films can become ferroelectric at 0 K outside the strain range of to (depending on Landau free energy coefficients), while considering the competing AFD phase. Although the strain fields here are not epitaxially induced, the strain tensor values, especially near dislocation cores and impurities (), exceed this threshold value. Although dislocations can result in other complex phenomena in STO, such as the interaction of dislocation cores with oxygen vacancies [70], dislocations have been shown to stabilize flexoelectric polarization at dislocation cores experimentally and theoretically [71]. Strain gradients are known to break local centrosymmetry, which can lead to the stabilization of a ferroelectric state in STO. If the magnitude of the strain field exceeds the threshold required for a ferroelectric state [23, 69], the strain fields induced by impurities, in the case of crystal B, could also stabilize local polarization.
@@
Another possibility for stabilization is the surface strain caused by the size of the microcrystals compared to the bulk. The atoms on the surface of the crystals experience a higher surface energy relative to the bulk. To relieve the surface energy, the atoms can distort themselves, thereby creating surface strain. Reconstructions of the microcrystals shown in Fig. 11, and figures 16, 17, and 20 in the appendices have regions of lattice expansion (yellow) on the surface, which lead to surface strains Fig. 6 and 14. The surface tension on these microcrystals could be analogous to the epitaxial strain induced on STO thin films [21], suggesting that the ferroelectric state of strained STO thin films [22, 23, 14] might also be relevant for these microcrystals.
@@
Although the ferroelectric polarization state has been proposed to be along the direction under compressive biaxial strain, and either along a or in-plane direction under tensile strain [23], the polarization state cannot be observed in the BCDI reconstructions. The ferroelectric phase will induce a polarization that changes the lattice parameter slightly, much like the AFD transition, but here we do not observe any split Bragg peaks. Ferroelectric polarization has been probed with BCDI [35], but the image of the ferroelectric state is not clear due to the dominant strain fields induced by the defects.
@@
We do not observe any noticeable changes in the heat maps (Figs. 4, 5, 12, and 13) corresponding to the cubic-to-tetragonal AFD phase transition at 105 K, nor do we observe a change in lattice parameter or change of slope (thermal expansion coefficient) in Fig. 15. Interestingly, He et al. reported that the AFD transition can occur as much as 50 K higher for biaxially strained STO thin films compared to bulk STO [22]. Their experiment measured a half-order superlattice Bragg peak associated with TiO6 octahedra rotation, and they observed a gradual decrease in intensity across different film thicknesses. Hence, it could be that the AFD phase transition occurs in crystal A and crystal B, but it cannot be detected from our measured crystal Bragg peaks or the lattice parameter change in Fig. 15. Furthermore, the rotation of the TiO6 octahedra around the crystallographic axis and the ratio for the tetragonal phase is below the strain resolution for BCDI. However, the AFD phase transition could also be influenced by the surface area-to-volume ratio and the chemistry of the microcrystals, which are not explored here.
@@ V Conclusion
This study successfully leverages the capability of MBCDI to resolve the residual strain and rotation tensor components for two STO microcrystals as they are cooled down to cryogenic temperatures. Below 50 K, each microcrystal transitioned to its low-temperature state, exhibiting increased strain magnitude and heterogeneity throughout the entire crystal. We do not observe the AFD phase transition in the MBCDI analysis, even though it should be present in strained samples [22]; however, it might be suppressed by a ferroelectric phase [12, 13]. Another possibility is the size or chemistry dependence of the hydrothermally synthesized STO that alters the phase transition behavior. Our results show structural modes in STO microcrystals at low temperatures, revealing information about structural transitions at the nanoscale.
@@ VI Data availability
The processed diffraction patterns, final reconstructions, and data analysis scripts are publicly available at [72].
@@ Acknowledgements.
We would like to thank Steve Shapiro for enlightening us about the neutron scattering work on STO, and Felix Hofmann for thoughtful discussions about dislocation analysis. Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. The authors acknowledge Diamond Light Source for time on Beamline I16 under proposal number MM30687, MM33417, and MM34617. N.Z. and Z.A. acknowledges support from the NSFC (Grant No. 12161141012) S. H., S. C., H. N., M. N., and H. K. acknowledges support from the National Research Foundation of Korea (Grant NRF-2021R1A3B1077076). Work performed at UCL was supported by EPSRC.Appendix A Phase retrieval
The reconstruction process for each CXDP was performed independently for each reflection. The cropped CXDP on the detector was padded with zeros to a size of pixels and binned by a factor of 2 along the detector plane, resulting in a data size of pixels.
Each reconstruction was initialized with a random guess. A guided phasing approach [73] with 60 (50 for the crystal shown in Fig. 11) individuals and 4 generations was used, employing a geometric average breeding mode. For each generation and population, a block of 20 error reduction (ER) and 180 hybrid input-output (HIO) iterations, with , was repeated three times. This was followed by 20 ER iterations to finalize the object. The shrinkwrap algorithm [74] with a threshold of 0.1 was used to update the real-space support at each iteration. The values for the Gaussian kernel shrinkwrap smoothing for each generation were , respectively. The best reconstruction was determined using a sharpness criterion, as it is an appropriate metric for crystals containing defects [75]. From the 30 (25 for the crystal shown in Fig. 11) best reconstructions, candidates well-correlated to the most similar candidate were averaged to produce the final reconstruction. This process was repeated again, with the reconstructions from the first round now as starting guesses.
The overall average 3D spatial resolution was 63.7 nm for crystal A and 61.0 nm for crystal B. This was determined by differentiating the electron density amplitude across the crystal/environment interface for five linearly independent directions and fitting a Gaussian to each of the profiles. The reported spatial resolution was the average full width at half maximum of the Gaussian profiles, across all measured reflections and temperatures for each crystal.
@@ Appendix B Individual reconstructions
The individual reconstructions for crystal A are shown in Fig. 16. Here, the phase vortices, shown within the purple circles, represented the positions of the dislocation cores [58] and were used to determine the dislocation lines discussed further in Appendix C. Fig. 17 shows the individual reconstructions for crystal B.